OpenStudy (anonymous):

Will medal Please help me Solve for t 100=60tan(pi/30(t+/-60))+/-60

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@DanJS

TheSmartOne (thesmartone):

Sorry I don't know :/

OpenStudy (kittiwitti1):

this is what i'm getting from your equation = v=;;\[100=60tan(\frac{\pi}{30}(t\pm60)\pm60)\]

OpenStudy (anonymous):

Yeah that's it

OpenStudy (anonymous):

The last 60 is outside the parenthesis

OpenStudy (kittiwitti1):

oh maybe that's why they were confused. but i suck with those signs so i can't help you sorry \[\pm \text{ < these signs i just don't get them}\]oh okay:\[100=60\tan(\frac{\pi}{30}(t\pm60))\pm60\]this?

OpenStudy (anonymous):

Yeah that's it

OpenStudy (kittiwitti1):

ok sorry i can't help you though u vu;;

OpenStudy (anonymous):

It's ok thanks

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

What have you done?

OpenStudy (anonymous):

Well when I tried to solve it I got 723 but I literally had no idea what I was doing or how I got that number

OpenStudy (anonymous):

I am completely lost

OpenStudy (anonymous):

I have no idea what you did for that either, but I guess we need two equations as it's ±

OpenStudy (anonymous):

\[\pm \]

OpenStudy (anonymous):

\[100=60\tan(\frac{\pi}{30}(t-60))-60 \implies 160 =60\tan(\frac{\pi}{30}(t-60))\] now divide by 60 on both sides \[\frac{ 8 }{ 3 } = \tan \left( \frac{ \pi }{ 30 }(t-60) \right)\] Now you can take the inverse \[\tan^{-1} \frac{ 8 }{ 3 } = \frac{ \pi }{ 30 }(t-60)\] \[t = 60 + \frac{ 30 \tan^{-1}(8/3) }{ \pi }\]

OpenStudy (anonymous):

Oh so at the end I'm going to have two answers

OpenStudy (anonymous):

Now do it with +

TheSmartOne (thesmartone):

Really interesting.... http://prntscr.com/5o8nc2

OpenStudy (anonymous):

You always seem to get weird problems huh :P, probably something was wrong with latex.